# Superfast Low Rank Approximation

### v2

@@ 5.1 Test overview
For both synthetic and real world input matrices the test results in our Table 5.1 show very good output accuracy of CUR LRA output by just five C-A loops (with uniform initialization), which was still quite good even where we applied the Cynical algorithm. The accuracy was significantly worse in the case of the Primitive algorithm. Tables 5.2 and 5.3 show the other test results for CUR LRA described at the end of Sec. 1.2.
We have performed the tests in the Graduate Center of the City University of New York by using MATLAB. Generating synthetic inputs we used Gaussian random matrices, filled with independent identically distributed normal (standard Gaussian) random variables. In particular we applied its standard normal distribution function “randn()” to generate Gaussian matrices and calculated $\epsilon$-ranks of matrices for $\epsilon=10^{-6}$ by using the MATLAB’s function ”rank(-,1e-6)”, which only counts singular values greater than $10^{-6}$. Some numerical experiments were executed with software custom programmed in $C^{++}$ and compiled with LAPACK version 3.6.0 libraries.
@@ 5.3 Four test algorithms
In our tests we applied and compared the following four algorithms for computing CUR LRA to input matrices $M$ having $\epsilon$-rank $r$:
  • Tests 1 (The Primitive algorithm for $k=l=r$): Uniformly choose two index sets $\mathcal{I}$ and $\mathcal{J}$, both of cardinality $r$, then compute a nucleus $U=M_{\mathcal{I},\mathcal{J}}^{-1}$ and define CUR LRA
@@
  • Tests 2 (Five loops of C–A): Uniformly choose an initial row index set $\mathcal{I}_{0}$ of cardinality $r$, then perform five loops of C–A (cf. Sec. 3.2) incorporating Alg. 1 of [50] as a subalgorithm that produces $r\times r$ CUR generators. At the end compute a nucleus $U$ and define CUR LRA as in Tests 1.
  • Tests 3 (A Cynical algorithm for $p=q=4r$ and $k=l=r$): Uniformly choose a row index set $\mathcal{K}$ and a column index set $\mathcal{L}$, both of cardinality $4r$, and then apply Algs. 1 and 2 from [50] to compute an $r\times r$ submatrix $M_{\mathcal{I},\mathcal{J}}$ of $M_{\mathcal{K},\mathcal{L}}$. Compute a nucleus and obtain CUR LRA by applying equation (5.1).
  • Tests 4 (Combination of a single C–A loop with Tests 3): Uniformly choose a column index set $\mathcal{L}$ of cardinality $4r$; then perform a single C–A loop (made up of a single horizontal step and a single vertical step): First by applying Alg. 1 from [50] define an index set $\mathcal{K}^{\prime}$ of cardinality $4r$ and the submatrix $M_{\mathcal{K}^{\prime},\mathcal{L}}$ in $M_{:,\mathcal{L}}$; then by applying this algorithm to matrix $M_{\mathcal{K}^{\prime},:}$ find an index set $\mathcal{L}^{\prime}$ of cardinality $4r$ and define submatrix $M_{\mathcal{K}^{\prime},\mathcal{L}^{\prime}}$ in $M_{\mathcal{K}^{\prime},:}$. Then proceed as in Tests 3 – find an $r\times r$ submatrix $M_{\mathcal{I},\mathcal{J}}$ in $M_{\mathcal{K}^{\prime},\mathcal{L}^{\prime}}$ by applying Algs. 1 and 2 from [50], compute a nucleus and CUR LRA.
@@ 5.5 LRA by means of random sampling and C-A acceleration
Whp the randomized algorithms of [20] run at superlinear cost and are expected to output an LRA within any fixed positive relative error norm bound $\epsilon$. We tested both the randomized algorithm of [20] and its combination with C-A iterations. Tables 5.2 and 5.3 display the relative errors $\frac{\|M-\tilde{M}\|}{\|M\|}$ of the LRA $\tilde{M}$ of the matrices $M$ computed in two ways: by means of [20, Alg. 2] (see the lines marked “CUR”) or eight C-A iterations with [20, Alg. 1] applied at all vertical and horizontal steps (see the lines marked “C-A”). The overall cost of performing the algorithms is superlinear in the former case and sublinear in the latter case. The columns of the tables marked with $\epsilon$-rank” display $\epsilon$-rank of an input matrix. The columns of the tables marked with ”k=lk=l” show the number of rows and columns in a square matrix of CUR generator.
As we pointed out at the end of Sec. 1.2, in almost all our tests this dramatic acceleration was achieved at the price of only minor deterioration of output accuracy.
@@
Table 5.2: LRA errors of Cross-Approximation (C-A) incorporating [20, Algorithm 1] in comparison to stand-alone CUR of [20, Algorithm 2].
@@
Table 5.3: LRA errors of Cross-Approximation (C-A) incorporating [20, Algorithm 1] in comparison to stand-alone CUR of [20, Algorithm 2].
@@ 6 LRA by means of sketching: four algorithms
Next we slightly generalize the sketching LRA algorithms of [39, 70] (see Remarks 6.16.3).
@@
We can obtain Algs. 6.2 by applying Alg. 6.1 to the transpose $M^{*}$. Likewise, by applying Alg. 6.3 to $M^{*}$ we obtain Alg. 6.4. We only study Algs. 6.1 and 6.3, but can readily extend that study to Algs. 6.2 and 6.4.
@@
Fix $k=r$, $l>k$, a random $n\times l$ matrix $H$ (e.g., Gaussian or SRHT matrix), identity matrix $S$, and $T$ equal to the $R$ factor in the $QR$ factorization of $MH$. Then the matrix $X$ has orthonormal columns and Algs. 6.1 turns into Proto-algorithm of [39, Sec. 9], while Alg. 6.3 turns into the Generalized Nyström algorithm of [48, Eqn. (3)]. [48] stabilizes the latter algorithm numerically – essentially by means of setting to 0 all singular values of the matrix $W$ exceeded by a fixed $\epsilon$, “a modest multiple of the unit roundoff $u$ times W||W||”. Our study can be readily extended to such a stabilized LRA because the stabilization little affects complexity of LRA and only improves its output accuracy. We can obtain various other modifications of Algs. 6.1 and 6.3 by fixing other sketching matrices $F$ and $H$.
@@
Column (resp. row) sketching turns into column (resp. row) subset selection where $H$ in Alg. 6.1 (resp. $F$ in Alg. 6.2) is a sampling matrix, that is, a full rank submatrix of a permutation matrix.
@@ 7.1 Deterministic error bounds of Range Finder
Next we recall some known estimates for the errors of Alg. 6.1, to be used in the next section.
@@
[39, Thm. 9.1]. Suppose that Alg. 6.1 has been applied to a matrix $M$ and let
@@
Under the assumptions of Thm. 7.1 and for $\tau_{r+1}(M)$ of Thm. 1.1 it holds that
@@ 7.2 Impact of pre-multiplication on the errors of LRA
The following theorem shows that the overall error bounds of Alg. 6.3 are dominated by the product of the norm $|W|||$ and the error norm bound of Alg. 6.1.
@@
See [70]. Let Alg. 6.3 output a matrix $XY$ for $Y=(FX)^{+}FM$ and let $m\geq k\geq l=\operatorname{rank}(X)$. Then
@@
if both matrices $F$ and $X$ have orthonormal columns. The sketch matrix $F$ is our choice, and for $X$ we ensure column orthogonality by properly choosing the matrix $T$ in Alg. 6.3 (see Remark 6.2).
@@ 8 Output error norm bounds for dual sketching algorithms
Given the matrices $MHT^{-1}$ and $S^{-1}FM$, Alg. 6.3 uses $O(kln)$ flops and hence runs superfast for $kl\ll m$. If also $l^{2}\ll m$ and $k^{2}\ll n$, then for proper Ultrasparse matrices $F$ and $H$ we can compute the matrices $MHT^{-1}$ and $S^{-1}FM$ superfast as well and hence can perform entire Algorithm 6.3 superfast, involving much less than $mn$ entries of $M$ and other scalars already where $k\ll m$ and $l\ll n$. Although such an LRA fails for a worst case input (cf. Appendix A), it succeeds whp in the case of any fixed, possibly sparse, well-conditioned matrix $H$ of full rank and a random input matrix $M$ that admits LRA, as we prove next. In view of the previous subsection we only need to prove this for Alg. 6.1.
@@ 8.1 Auxiliary results
Constant matrices are filled with constants, unlike random matrices, filled with random variables.
Theorem 8.1.
Proof.
Rank deficiency of matrices $H$, FMFM, and $MH$ is equivalent to turning into 0 the determinants $\det(FF^{*})$, $\det(H^{*}H)$, $\det((MH^{*})MH)$, and $\det(FM(FM)^{*})$, respectively. The claim follows because these equations define algebraic varieties of lower dimension in the linear spaces of the entries, considered independent variables (cf., e.g., [7, Prop. 1]). ∎
Remark 8.1.
Events that occur with probability 0 are immaterial for our probability estimates, and hereafter we say that a matrix has full rank even if it is rank deficient with probability 0.
Lemma 8.2.
[Orthogonal Invariance.] [68, Theorem 3.2.1]. Suppose that $G$ is an $m\times n$ Gaussian matrix, $k\leq\min\{m,n\}$ is a positive integer, and $S\in\mathbb{R}^{k\times m}$ and $T\in\mathbb{R}^{n\times k}$ are constant matrices, having orthonormal rows and columns, respectively. Then $SG$ and $GT$ are random matrices having distribution of $k\times n$ and $m\times k$ Gaussian random matrices, respectively.
Hereafter write $e:=2.7182818\dots$.
@@
The matrix $GH$ has distribution of an $r\times l$ Gaussian random matrix by virtue of Lemma 8.2, and hence we can assume that it has full rank (see Remark 8.1). Now recall Thm. C.1 and claim (ii) of Thm. C.2 and obtain
@@
We can extend our lower bounds of Lemma 8.3 and Cor. 8.1 on $\sigma_{r}(QH)$ to the case of any matrix $H$ of full rank $l$ if we decrease these bounds by a factor of $\sigma_{l}(H)$.
8.2 Output errors of Alg. 6.1 for a matrix with random singular space
@@
[Errors of Alg. 6.1 for an input with a random singular space.] Suppose that $G$ is an $r\times n$ Gaussian matrix, $H\in\mathbb{R}^{n\times l}$ is a constant matrix, $n>36r$, l>22(r1)l>22(r-1), $r\leq l<\min(m,n)$, and Alg. 6.1 applied to the matrix $M$ of (8.1) outputs two matrices $X$ and $Y$.
@@
We can assume that the matrices $G$ and $GH$ have full rank $r$ (see Thm. 8.1 and Remark 8.1). Consider SVD
@@
write $C_{1}:=Q^{*}H$, apply Cor. 7.1, and deduce that
@@
(i) Recall from Cor. 8.1 that
@@
Combine Thm. 7.1, equation $Q_{\perp}||_{2}=1$, and the bound $\sigma_{r}(C_{1})\geq\frac{1}{4}\sqrt{l/n}\cdot\sigma_{r}(H)$ and obtain

### v3

@@ 6.1 Test overview
For both synthetic and real world input matrices the test results in our Table 6.1 show good accuracy of CUR LRA output by just five C-A loops (with uniform initialization). Accuracy was still quite good where we applied the Cynical algorithm but significantly worse where we applied the Primitive algorithm. Tables 6.2 and 6.3 compare the output accuracy of the random sampling algorithms of [16] and their superfast combination with C-A iterations.
We have performed the tests in the Graduate Center of the City University of New York by applying MATLAB. We generated synthetic inputs by using Gaussian matrices, which we created by using standard normal distribution function “randn()”. We calculated $\epsilon$-ranks of matrices for $\epsilon=10^{-6}$ by using the MATLAB’s function ”rank(-,1e-6)”, which only counts singular values greater than $10^{-6}$. Some numerical experiments were executed with software custom programmed in $C^{++}$ and compiled with LAPACK version 3.6.0 libraries.
@@ 6.3 Four test algorithms
We tested the following four algorithms for computing CUR LRA of input matrices $M$ having $\epsilon$-rank $r$:
  • Tests 1 (The Primitive algorithm for $k=l=r$): Uniformly sample two index sets $\mathcal{I}$ and $\mathcal{J}$, both of cardinality $r$, then compute a nucleus $U=M_{\mathcal{I},\mathcal{J}}^{-1}$ and define CUR LRA
@@
  • Tests 2 (Five loops of C–A): Uniformly sample an initial row index set $\mathcal{I}_{0}$ of cardinality $r$, then perform five loops of C–A (cf. Sec. 3.2) incorporating Alg. 1 of [41] as a subalgorithm that produces $r\times r$ CUR generators. At the end compute a nucleus $U$ and define CUR LRA as in Tests 1.
  • Tests 3 (A Cynical algorithm for $p=q=4r$ and $k=l=r$): Uniformly sample a row index set $\mathcal{K}$ and a column index set $\mathcal{L}$, both of cardinality $4r$. Then apply Algs. 1 and 2 of [41] to compute an $r\times r$ submatrix $M_{\mathcal{I},\mathcal{J}}$ of $M_{\mathcal{K},\mathcal{L}}$. Compute a nucleus and obtain CUR LRA of (6.1).
  • Tests 4 (Combination of a single C–A loop with Tests 3): Uniformly sample a column index set $\mathcal{L}$ of cardinality $4r$; then perform a single C–A loop: first define an index set $\mathcal{K}^{\prime}$ of cardinality $4r$ and the submatrix $M_{\mathcal{K}^{\prime},\mathcal{L}}$ in $M_{:,\mathcal{L}}$ by applying Alg. 1 of [41]; then apply this algorithm to the matrix $M_{\mathcal{K}^{\prime},:}$ to find an index set $\mathcal{L}^{\prime}$ of cardinality $4r$ and to define submatrix $M_{\mathcal{K}^{\prime},\mathcal{L}^{\prime}}$ in $M_{\mathcal{K}^{\prime},:}$. Then proceed as in Tests 3 – find an $r\times r$ CUR generator $M_{\mathcal{I},\mathcal{J}}$ in $M_{\mathcal{K}^{\prime},\mathcal{L}^{\prime}}$ by applying Algs. 1 and 2 of [41] and finally compute a nucleus and CUR LRA of $M$ based on (2.3) for $G:=M_{\mathcal{I},\mathcal{J}}$.
@@ 6.5 LRA by means of random sampling and C-A acceleration
Tables 6.2 and 6.3 display the relative errors $\frac{\|M-\tilde{M}\|}{\|M\|}$ of the LRA $\tilde{M}$ of the matrices $M$ computed in two ways: by means of [16, Alg. 2] (see the lines marked “CUR”), which are expected to be near-optimal under the Frobenius morm whp, or eight C-A iterations combined with [16, Alg. 1] (see the lines marked “C-A”). The algorithms run at superlinear cost in the former case and are superfast in the latter case. The columns of the tables marked with $\epsilon$-rank” display $\epsilon$-rank of an input matrix. The columns of the tables marked with ”k=lk=l” show the number of rows and columns in a square matrix of CUR generator.
As we pointed out in Sec. 1.2.1, in almost all our tests this dramatic acceleration was achieved at the price of only minor deterioration of output accuracy.
@@
Table 6.2: LRA errors of Cross-Approximation (C-A) incorporating [16, Algorithm 1] in comparison to stand-alone CUR of [16, Algorithm 2].
@@
Table 6.3: LRA errors of Cross-Approximation (C-A) incorporating [16, Algorithm 1] in comparison to stand-alone CUR of [16, Algorithm 2].
@@ 7 LRA by means of sketching: four algorithms
Next we slightly generalize the sketching LRA algorithms of [30, 60] (see Remarks 7.17.3); Thms. 8.2 and 9.1 combined formally support output accuracy of dual version of all of them whp; we applied our tests to the algorithms of [30, 60] accelerated by means of using Abridged SRHT multipliers $F$ and $H$.
@@
We can obtain Algs. 7.2 by applying Alg. 7.1 to the transpose $M^{*}$. Likewise, by applying Alg. 7.3 to $M^{*}$ we obtain Alg. 6.4. We only study Algs. 7.1 and 7.3, but can readily extend that study to Algs. 7.2 and 6.4.
@@
Fix $k=r$, $l>k$, a random $n\times l$ matrix $H$ (e.g., Gaussian or SRHT matrix), identity matrix $S$, and $T$ equal to the $R$ factor in the $QR$ factorization of $MH$. Then the matrix $X$ has orthonormal columns and Algs. 7.1 turns into Proto-algorithm of [30, Sec. 9], while Alg. 7.3 turns into the Generalized Nyström algorithm of [39, Eqn. (3)]. [39] stabilizes the latter algorithm numerically – essentially by means of setting to 0 all singular values of the matrix $W$ exceeded by a fixed $\epsilon$, “a modest multiple of the unit roundoff $u$ times W||W||”. Our study can be readily extended to such a stabilized LRA because the stabilization little affects complexity of LRA and only improves its output accuracy. We can obtain various other modifications of Algs. 7.1 and 7.3 by fixing other sketching matrices $F$ and $H$.
@@
Column (resp. row) sketching turns into column (resp. row) subset selection where $H$ in Alg. 7.1 (resp. $F$ in Alg. 7.2) is a sampling matrix, that is, a full rank submatrix of a permutation matrix.
@@ 8.1 Deterministic error bounds of Range Finder
Next we recall some known estimates for the errors of Alg. 7.1, to be used in the next section.
@@
[30, Thm. 9.1]. Suppose that Alg. 7.1 has been applied to a matrix $M$ and let
@@
Under the assumptions of Thm. 8.1 and for $\tau_{r+1}(M)$ of Thm. 1.1 it holds that
@@ 8.2 Impact of pre-multiplication on the errors of LRA
The following theorem shows that the overall error bounds of Alg. 7.3 are dominated by the product of the norm $|W|||$ and the error norm bound of Alg. 7.1.
@@
See [60]. Let Alg. 7.3 output a matrix $XY$ for $Y=(FX)^{+}FM$ and let $m\geq k\geq l=\operatorname{rank}(X)$. Then
@@
if both matrices $F$ and $X$ have orthonormal columns. The sketch matrix $F$ is our choice, and for $X$ we ensure column orthogonality by properly choosing the matrix $T$ in Alg. 7.3 (see Remark 7.2).
@@ 9 Output error norm bounds for dual sketching algorithms
Given the matrices $MHT^{-1}$ and $S^{-1}FM$, Alg. 7.3 uses $O(kln)$ flops and hence runs superfast for $kl\ll m$. If also $l^{2}\ll m$ and $k^{2}\ll n$, then for proper Ultrasparse matrices $F$ and $H$ we can compute the matrices $MHT^{-1}$ and $S^{-1}FM$ superfast as well and hence can perform entire Algorithm 7.3 superfast, involving much less than $mn$ entries of $M$ and other scalars, already where $k\ll m$ and $l\ll n$. Although such an LRA fails for a worst case input, it succeeds whp under our dual model in the case of any fixed, possibly sparse, well-conditioned multipliers $F$ and $H$ of full rank and a random input matrix $M$ that admits LRA. In view of the previous subsection we only need to prove this for Alg. 7.1.
@@ 9.1 Auxiliary results
The matrix $GH$ has distribution of an $r\times l$ Gaussian matrix by virtue of Lemma 1.3, and hence we can assume that it has full rank (see Remark 1.2). Now recall Thm. C.1 and claim (ii) of Thm. C.2 and obtain
@@
We can extend our lower bounds of Lemma 9.2 and Cor. 9.1 on $\sigma_{r}(QH)$ to the case of any matrix $H$ of full rank $l$ if we decrease these bounds by a factor of $\sigma_{l}(H)$.
9.2 Output errors of Alg. 7.1 for a matrix with random singular space
@@
[Errors of Alg. 7.1 for an input with a random singular space.] Suppose that $G$ is an $r\times n$ Gaussian matrix, $H\in\mathbb{R}^{n\times l}$ is a constant matrix, $n>36r$, l>22(r1)l>22(r-1), $r\leq l<\min(m,n)$, and Alg. 7.1 applied to the matrix $M$ of (9.1) outputs two matrices $X$ and $Y$.
@@
We can assume that the matrices $G$ and $GH$ have full rank $r$ (see Thm. 1.3 and Remark 1.2). Consider SVD
@@
write $C_{1}:=Q^{*}H$, apply Cor. 8.1, and deduce that
@@
(i) Recall from Cor. 9.1 that
@@
Combine Thm. 8.1, equation $Q_{\perp}||_{2}=1$, and the bound $\sigma_{r}(C_{1})\geq\frac{1}{4}\sqrt{l/n}\cdot\sigma_{r}(H)$ and obtain
@@
Bound the output errors of Algs. 7.3 and 3.4 by combining the estimates of this section and Sec. 8.2 and by transposing the input matrix $M$.
9.3 Output errors of Range Finder for a perturbed factor-Gaussian input
Assumption 9.2.
For an $r\times n$ Gaussian matrix $G$ and a constant matrix $A\in\mathbb{R}^{m\times r}$ of full rank $r<\min(m,n)$ define the matrices $B:=\frac{1}{\sqrt{n}}\cdot G$ and $\tilde{M}:=AB$ and call $M:=\tilde{M}+E$ a perturbed right factor-Gaussian matrix if the Frobenius norm of a perturbation matrix $E$ is sufficiently small in comparison to $\sigma_{r}(A)$, as we specify in the next theorem.
Theorem 9.2.
[Errors of Range Finder for a perturbed right factor-Gaussian matrix.] Suppose that we are given an $r\times n$ Gaussian matrix $G$ and constant matrices $H\in\mathbb{R}^{n\times l}$, $A\in\mathbb{R}^{m\times r}$, and $E\in\mathbb{R}^{m\times n}$ for $r\leq l<\min(m,n)$. Let $\tilde{M}$ be a right factor-Gaussian matrix of Assumption 9.2 and let $M=\tilde{M}+E$ for a perturbation matrix $E$. Apply Alg. 7.1 to the matrix $M$ with a test matrix $H$, and let $X$ and $Y$ denote the output matrices. Assume that $n>36r$ and l>22(r1)l>22(r-1).
(i) Suppose that the matrix $H$ has orthonormal columns and that $E||_{F}\leq\frac{\sigma_{r}(A)}{48\sqrt{n/l}+6}$. Then
$|M-XY|||/\tilde{\sigma}_{r+1}(M)\leq\sqrt{1+100n/l}$
with a probability no less than $1-\exp(-\frac{l-r}{20})-\exp(-\frac{n-r}{20})-\exp(-\frac{n}{72})$.
[…]
$|M-XY|||/\tilde{\sigma}_{r+1}(M)\leq\sqrt{1+100\kappa^{2}(H)n/l}$
with a probability no less than $1-\exp(-\frac{l-r}{20})-\exp(-\frac{n-r}{20})-\exp(-\frac{n}{72}).$
Proof.
Recall that the matrices $B$, ABAB, and $BH$ have full rank with probability 1 (see Thm. 1.3) and assume that they do have full rank (see Remark 1.2). Let
$M=\begin{bmatrix}U_{r}&U_{\perp}\end{bmatrix}\begin{bmatrix}\Sigma_{r}&0\\ 0&\Sigma_{\perp}\end{bmatrix}\begin{bmatrix}V_{r}^{T}\\ V_{\perp}^{T}\end{bmatrix}\textrm{ and }\tilde{M}=\begin{bmatrix}\tilde{U}_{r}&\tilde{U}_{\perp}\end{bmatrix}\begin{bmatrix}\tilde{\Sigma}_{r}&0\\ 0&0\end{bmatrix}\begin{bmatrix}\tilde{V}_{r}^{T}\\ \tilde{V}_{\perp}^{T}\end{bmatrix}$
be SVDs, where $V_{r}$ and $\tilde{V}_{r}$ are the matrices of the $r$ top right singular vectors of $M$ and $\tilde{M}$, respectively. Define $C_{1}=V_{r}^{T}H$ and $\tilde{C}_{1}:=\tilde{V}_{r}^{T}H$ as in (8.1). Now Thm. 8.1, that is, [30, Thm. 9.1], implies that
[…]
Recall that the columns of $V_{r}$ span the row space of an $r\times n$ Gaussian matrix $G$ and deduce from Corollary 9.1 that $\sigma_{r}(\tilde{C}_{1})>\frac{1}{4}\sqrt{l/n}$ with a probability no less than $1-\exp(-\frac{l-r}{20})-\exp(-n/72)$. It remains to estimate how the perturbation $E$ alters the leading right singular space of $\tilde{M}$ and how close is $\sigma_{r}(C_{1})$ to $\sigma_{r}(\tilde{C}_{1})$.
[…]
Claim (ii) of Thm. C.2 implies that whp the $r$-th singular value of $\tilde{M}=AB$ is not much less than the $r$-th singular value of $A$. Readily deduce from Cor. 9.1 that
$\operatorname{Probability}\{\sigma_{r}(\tilde{M})<\sigma_{r}(A)/3\}\leq\operatorname{Probability}\{\sigma_{r}(B)=\frac{1}{\sqrt{n}}\sigma_{r}(G)<1/3\}\leq e^{-(n-r)/20}.$
Hence $E||_{F}\leq\frac{\sigma_{r}(A)}{48\sqrt{n/l}+6}\leq\frac{\sigma_{r}(\tilde{M})}{16\sqrt{n/l}+2}$ with a probability no less than $1-e^{-(n-r)/20}$, and so $P||_{2}\leq\frac{1}{8}\sqrt{l/n}$ for some matrix $P$ of Lemma 1.2 such that $P||_{2}\leq\frac{1}{8}\sqrt{l/n}$.
Now let this holds, let $\sigma_{r}(\tilde{C}_{1})>\frac{1}{4}\sqrt{l/n}$, and deduce that
$\sigma_{r}(V_{r}^{T}H)$ $=\sigma_{r}\big{(}(I_{r}+P^{T}P)^{-1/2}(\tilde{V}_{r}^{T}+P^{T}\tilde{V}_{\perp}^{T})H\big{)}$ (9.3)
$\geq\sigma_{r}\big{(}(I_{r}+P^{T}P)^{-1/2}\big{)}\sigma_{r}(\tilde{V}_{r}^{T}H+P^{T}\tilde{V}_{\perp}^{T}H)$ (9.4)
$\geq\frac{1}{\sqrt{1+(\sigma_{1}(P))^{2}}}\big{(}\sigma_{r}(\tilde{C}_{1})-\sigma_{1}(P)\big{)}>\frac{1}{10}\sqrt{l/n}.$ (9.5)
Equality (9.3) holds because the matrix $\tilde{V}_{r}+\tilde{V}_{\perp}P$ is normalized by $(I_{r}+P^{T}P)^{-1/2}$ (see Remark 1.1) and has the same column span as $V_{r}$. By applying the union bound deduce that inequality (9.5) holds with a probability no less than $1-\exp(-\frac{l-r}{20})-\exp(-\frac{n-r}{20})-\exp(-\frac{n}{72})$.
To prove claim (ii), we essentially need to show that $\sigma_{r}(C_{1})=\sigma_{r}(V_{r}^{T}H)\geq\frac{1}{10}\sqrt{l/n}\cdot\sigma_{l}(H)$, and then the claim will follow readily from inequality (9.3). Let $H=U_{H}\Sigma_{H}V_{H}^{T}$ be compact SVD, such that $U_{H}\in\mathbb{R}^{n\times l}$ , $\Sigma_{H}\in\mathbb{R}^{l\times l}$, and $V_{H}\in\mathbb{R}^{l\times l}$, and obtain that $\sigma_{r}(\tilde{C}_{1})\geq\sigma_{r}(\tilde{V}_{r}^{T}U_{H})\sigma_{l}(\Sigma_{H})$ and
$\operatorname{Probability}\{\sigma_{r}(\tilde{C}_{1})<\frac{1}{4}\sqrt{l/n}\cdot\sigma_{l}(H)\}<\exp(-\frac{l-r}{20})+\exp(-\frac{n}{72}).$
Next bound $\sigma_{r}(C_{1})$ by showing that the column spaces of $V_{r}$ and $\tilde{V}_{r}$ are sufficiently close to one another if the perturbation $V_{r}-\tilde{V}_{r}$ is sufficiently small. Namely, assume that $E||_{F}\leq\frac{\sigma_{r}(A)}{12}$, and then the assumptions of Lemma 1.2 holds whp. By applying the same argument as in the proof of claim (i), deduce that
[…]
with a probability no less than $1-e^{-(n-r)/20}$. It follows that $P||_{2}\leq\frac{1}{8}\sqrt{l/n}\cdot\sigma_{l}(H)$ for some matrix $P$ of Lemma 1.2. Hence $\sigma_{r}(C_{1})\geq\frac{1}{10}\sqrt{l/n}\cdot\sigma_{l}(H)$ whp. ∎